OpenStudy (anonymous):

When 0.236 g of the hydrate CoCl2•xH2O is heated, 0.128 grams of anhydrous CoCl2 remains. What is the value of x in the formula for the hydrate? (hint: remember that x represents the ratio of moles of H2O to moles of hydrate compound, and that moles of hydrate compound equals the moles of anhydrous compound)

OpenStudy (frostbite):

You do as they suggest in the hint however I help you a little more on the right track: First We are interested in evaluating the expression: \[\Large x=\frac{ n(\sf H_2O )}{ n( \sf CoCl_2) }\] By by using that \[\Large n(H_2O)=\frac{ m(H_2O) }{ M(H_2O }\]\[\Large n(CoCl_2)= \frac{ m(CoCl_2) }{ M(CoCl_2) }\] We can calculate the amount of substance for each of the compounds. The mass of water water in the hydrate must be the amount of anhydrous minus hydrate giving us the expression: \[\large m(H_2O)=(0.236-0.128)~g=0.108~g\] The molar mass \(M\) can be calculated using your periodic table as the sum of all the atoms in the molecule. When calculating \(x\) round up or down to the nearest whole number. From here it should be straight forward. If any questions feel free writing.